OpenStudy (anonymous):

simplify expression e^-2-f^-1/ef

OpenStudy (ddcamp):

Try entering that in the equation editor, it's a bit unclear.

OpenStudy (abb0t):

Can you rewrite it using LaTex? That's a bit difficult to read. It can be taken various ways.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ e ^{-2}-f ^{-1} }{ ef }\]

OpenStudy (abb0t):

\(\sf e^{-2}=\frac{1}{e^2}\), and similarly, \(\sf f^{-1}=\frac{1}{f}\)

OpenStudy (anonymous):

@Sephora f-e^2/f^2e^3

OpenStudy (anonymous):

tequila i dont understand

OpenStudy (anonymous):

@Sephora That's the answer

OpenStudy (anonymous):

how did you got that answer please show me. I want to learn

OpenStudy (abb0t):

Don't just give answers @tequila4 that's not very helpful for the user and as you can see, they are trying to learn.

OpenStudy (anonymous):

Thank you abbot

OpenStudy (abb0t):

It might help if you separated them into two fractions: \(\sf \color{red}{\frac{e^{-2}}{ef}-\frac{f^{-1}}{ef}}\)

OpenStudy (anonymous):

okay

OpenStudy (abb0t):

Remember, when you have a negative exponent like you do here: \(\sf \color{blue}{e^{-2} ~it~simply~means:\frac{1}{e^2}}\)

OpenStudy (abb0t):

So, you actually have: \(\sf \color{red}{\frac{1}{e^2(ef)}-\frac{1}{f(ef)}}\)

OpenStudy (abb0t):

Do you follow me so far, @Sephora ?

OpenStudy (anonymous):

yes

OpenStudy (abb0t):

You then multiply across. Remember that when you multiply across parenthesis, you are actually ADDING their exponents, so it's \(\sf \color{red}{e(ef) = e^{1+1}f = e^2f}\), now you try the second one.

OpenStudy (hesan):